OpenStudy (jamanorthi):

insert grouping symbole in the expression so that the value of the expression is 14 1. 9+39+22/11-9+3

OpenStudy (petiteme):

There xD

OpenStudy (jamanorthi):

@petiteme yeah

OpenStudy (petiteme):

\[\frac{ 9+39+32 }{ 11-9+3}\] Since the numerator is all addition, you can just add it all. So it'll be \[\frac{ 70 }{ 11-9+3 }\] Look at the denominator. What if you group it like this: 11-(9+3) Will you get 12? Or if you group it like this: (11-9) + 3

OpenStudy (jamanorthi):

okay let me solve it now

OpenStudy (petiteme):

Okay. I'll wait ^_^

OpenStudy (jamanorthi):

@petiteme i don't get it

OpenStudy (petiteme):

Okay lemme help you out. To get 14 as an answer, try solving this first. \[\frac{ 70 }{ (11-9) + 3 }\] That's out first trial of grouping.

OpenStudy (jamanorthi):

okay

OpenStudy (jamanorthi):

70/23

OpenStudy (petiteme):

Wrong. The first thing to do is perform the operation inside the parenthesis :)

OpenStudy (jamanorthi):

okay

OpenStudy (petiteme):

I'll wait ^_^

OpenStudy (jamanorthi):

i am so confused

OpenStudy (petiteme):

Okay. Do (11-9) and then when you get the answer, add it to 3.

OpenStudy (jamanorthi):

5

OpenStudy (petiteme):

Good! Then 5 will be your denominator right? \[\frac{ 70 }{ 5 }\] So 70 divided by 5 is 14 right?

OpenStudy (jamanorthi):

yeah it is

OpenStudy (petiteme):

So the grouping that you need is \[\frac{ (9+ 32) + 22 }{ (11-9) + 3 }\]

OpenStudy (jamanorthi):

okay thanks

OpenStudy (petiteme):

You're welcome :) I hope I helped you out.

OpenStudy (jamanorthi):

but it doesn't give 14 at the end when u solve it